Cheyenne Stores Inc., which uses the conventional retail inventory method, wishes to change to the LIFO retail method beginning with the accounting year ending December 31, 2020.

Amounts as shown below appear on the store’s books before adjustment.

Cost

Retail

Inventory, January 1, 2020 $14,537 $25,200
Purchases in 2020 116,088 199,800
Markups in 2020 12,500
Markdowns in 2020 5,000
Sales revenue in 2020 189,700


You are to assume that all markups and markdowns apply to 2020 purchases, and that it is appropriate to treat the entire inventory as a single department.

Compute the inventory at December 31, 2020, under the following methods.

A. The conventional retail method.

Ending inventory using the conventional retail method

B. The last-in, first-out retail method, effecting the change in method as of January 1, 2020. Assume that the cost-to-retail percentage for 2019 was recomputed correctly in accordance with procedures necessary to change to LIFO. This ratio was 62%.

Ending inventory using the last-in, first-out retail method
Answer #1

Solution A :

cost of goods available for sale:

Cost Retail
Inventory, January 1, 2020
$14,537
$25,200
Purchases in 2020 $116,088 $199,800
Markups in 2020 $12,500
Cost of goods available for sale $130,617 $237,500

cost to retail ratio = 130617/237500 = 0.55

calculation of ending inventory:

Cost Retail
Cost of Goods Available for Sale $130,617

$237,500

- mark down $5,000
− Sales $189,700
Ending Inventory $42,800
× Cost to Retail Ratio 0..55
Ending Inventory $23,540

Solution B :

cost of goods available for sale:

Cost Retail
Inventory, January 1, 2020
$14,537
$25,200
Purchases in 2020 $116,088 $199,800
Markups in 2020 $12,500
Mark down in 2020 $5,000
purchases available for sale $116,088 $207,300
Cost of goods available for sale $130,617 $232,500

cost to retail ratio = 116088/207300 = 0.56

calculation of ending inventory:

Cost Retail
Cost of Goods Available for Sale $130,617

$237,500

− Sales $189,700
Ending Inventory $47,800
Ending Inventory Cost Retail
Inventory, January 1, 2020 $14,537 $25,200
Purchases in 2020[(47800-25200)*0.56] 12656
27193